40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Mar 19 15:34:41 2020, anchal, DailyProgress, North Cavity, North path's buggy nature solved IMG_20200319_154013.jpg
    Reply  Sun Mar 22 17:23:29 2020, anchal, DailyProgress, North Cavity, North path's buggy nature NOT solved 
       Reply  Wed Mar 25 15:50:57 2020, anchal, DailyProgress, North Cavity, North path's buggy nature NOT solved Screen_Shot_2020-03-25_at_3.48.24_PM.png
          Reply  Tue Apr 14 12:24:16 2020, anchal, DailyProgress, North Cavity, North path's buggy nature Strikes Again! NorthPathWentBuggy.pdf
             Reply  Thu Apr 16 18:08:56 2020, anchal, DailyProgress, North Cavity, North path stopped being buggy NorthPathStoppedBeingBuggy.pdf
Message ID: 2564     Entry time: Sun Mar 22 17:23:29 2020     In reply to: 2563     Reply to this: 2565
Author: anchal 
Type: DailyProgress 
Category: North Cavity 
Subject: North path's buggy nature NOT solved 

It seems like this was never properly solved. On Friday, the same problem was back again. After trying relocking PMC and FSS on the north path without any luck, I switched off the laser to standby mode and after a minute restarted it and the problem went away. I have a strong suspicion that this problem has something to do with the laser temperature controller on the laser head itself. During the unstable state, I see a spike that starts a large surge in error signal of FSS loop which occurs every 1 second! (so something at 1 Hz). The loop actually kills the spike successfully withing 600-700 ms but then it comes back again. I'm not sure what's wrong, but if this goes on and the lockdown is enforced due to Corona virus, I won't even able to observe the experiment from a distance :(. I have no idea what went wrong here.

ELOG V3.1.3-